OpenStudy (anonymous):

Use the given values of x to graph each function. Use the graph to determine the domain and range of the function. 1. ; x = -3, -2, -1, 0, 1, 2, 3 Ok I got 6 similar probs so who wants to show me how to do them?

OpenStudy (anonymous):

@jim_thompson5910 @kirbykirby

OpenStudy (anonymous):

plz

OpenStudy (anonymous):

f(x)=x^2+1; x=-3,-2,-1,0,1,2,3 I need to graph it and then determine the range and domain

OpenStudy (anonymous):

where do i start?

jimthompson5910 (jim_thompson5910):

if x = -3, then f(x) = x^2 + 1 f(-3) = (-3)^2 + 1 f(-3) = ??

OpenStudy (anonymous):

f(-3)=10

OpenStudy (anonymous):

ok so how do I graph that?

jimthompson5910 (jim_thompson5910):

this means that the point (-3, 10) is on the graph

jimthompson5910 (jim_thompson5910):

when x = -2, f(-2) = ??

OpenStudy (anonymous):

ooooooooooo

jimthompson5910 (jim_thompson5910):

yeah you're just generating a bunch of points, plotting them, then drawing a curve through them

OpenStudy (anonymous):

coolio u r awesome!!!!!!!!!!!!!!!!!!!!!!!!!!!

jimthompson5910 (jim_thompson5910):

the more points, the more accurate the graph is, glad it's clicking now

OpenStudy (anonymous):

yes very much ty so much!!!!!!!!!!!!

jimthompson5910 (jim_thompson5910):

yw

OpenStudy (anonymous):

wait @jim_thompson5910 how do I find the domain and range?

jimthompson5910 (jim_thompson5910):

can you think of any number you can plug in for x, but it won't be defined for f(x)?

OpenStudy (anonymous):

no

jimthompson5910 (jim_thompson5910):

since any real number works for x, and it produces some number f(x), this means that the domain is the set of all real numbers

OpenStudy (anonymous):

o, what about the range?

jimthompson5910 (jim_thompson5910):

did you graph x^2 + 1

OpenStudy (anonymous):

no

jimthompson5910 (jim_thompson5910):

go ahead and do so

OpenStudy (anonymous):

so graph all 7 points?

jimthompson5910 (jim_thompson5910):

yes, then draw a curve through all of them

OpenStudy (anonymous):

ok just a minute

jimthompson5910 (jim_thompson5910):

ok

OpenStudy (anonymous):

Ok so I drew them and now i have a curve

jimthompson5910 (jim_thompson5910):

what is the lowest point

jimthompson5910 (jim_thompson5910):

you should have graphed a U shaped curve

OpenStudy (anonymous):

yup, 0,1 is the lowest point

jimthompson5910 (jim_thompson5910):

so the range is the set of all y values greater than or equal to 1 since y = 1 is the smallest y value possible

OpenStudy (anonymous):

so range=1?

jimthompson5910 (jim_thompson5910):

no the range is a set of numbers, not just 1 (it includes 1 though)

jimthompson5910 (jim_thompson5910):

y = 2, y = 3, ... are part of the range as well

OpenStudy (anonymous):

o range = =>1?

jimthompson5910 (jim_thompson5910):

y >= 1 yes

OpenStudy (anonymous):

ok and y=range

jimthompson5910 (jim_thompson5910):

that basically tells me that any value larger (or equal to) 1 will work for the range

jimthompson5910 (jim_thompson5910):

yeah

jimthompson5910 (jim_thompson5910):

y values on the graph y values for the points on the graph

OpenStudy (anonymous):

:) ty!!!!!!!!

jimthompson5910 (jim_thompson5910):

np

OpenStudy (anonymous):

u r awesome!!!!!!!!!!!!!!!!!!!!!!

jimthompson5910 (jim_thompson5910):

thanks